What is the total interest made after 4 years on a simple interest loan that accumulates 13% each year and has original amount was $1350.


1
Expert's answer
2021-06-28T16:44:33-0400

"SI=P\times R\times T\\=\$1350\times\frac{13}{100}\times4"

"=\$702"


the total interest made after 4 years is $702